# Ziliang Zhang College of Oceanic and Atmospheric Sciences

### v2

@@ 3.1 Performance Evaluation on ERA5 Dataset
During the 59 time steps (14.75 days) of the forecasting process, the model’s mean absolute error(MAE) is 0.08 m, with a minimum ACC of 0.89 and a maximum RMSE of 0.02 m. Considering the evaluation curves, the model performs well in forecasting the wave field. It is worth noting that although the model’s error decreases over time, it does not continue to decline like other forecasting models. Instead, it stabilizes at around 0.89 after the initial decrease. Based on this, we hypothesize that the model has learned the physical properties of waves.We will present relevant experiments in Section 3.3.
@@
During the 35-time-step (8.75 days) forecasting period, within the project area (longitude 120°E to 170°E, latitude 0°N to 50°N), the MAE was 0.24 m, the minimum ACC was 0.98, and the maximum RMSE was 0.04 m. Considering the evaluation curves, the model’s forecasting performance during Typhoon Goni can be regarded as satisfactory.
@@
During the 23-time-step (5.75 days) forecasting period, within the forecasting area (longitude 100°W to 60°W, latitude 0°N to 20°N), the MAE was 0.21 m, the minimum ACC was 0.84, and the maximum RMSE was 0.05 m. Considering the evaluation curves, the model’s forecasting performance in the hurricane region is also highly satisfactory.
@@ 3.2 Performance Comparison against Traditional Forecasting Models
During the first 7 time steps (1.75 days) of the forecasting process, the MAE of the OCN model was 0.14, which was lower than the MAE of the WW3 model (0.30). In terms of ACC, the minimum value for the OCN model was 0.99, higher than the 0.98 of the WW3 model. Regarding RMSE, the maximum value for the OCN model was 0.02, lower than the 0.03 of the WW3 model. Considering the evaluation curves, we can conclude that the short-term forecasting performance of the OCN model under normal conditions is better than that of the WW3 model.
@@
During the 11-time-step (2.75 days) forecasting period, within the forecasting area (longitude 120°E to 162°E, latitude 0°N to 30°N), the MAE of the OCN model (0.16 m) was lower than that of the WW3 model (0.26 m). In terms of ACC, the minimum value for the OCN model (0.98) was higher than that of the WW3 model (0.97). Regarding RMSE, the maximum value for the OCN model (0.07 m) was lower than that of the WW3 model (0.09 m). Considering the evaluation curves, we can conclude that the OCN model performed better than the WW3 model in forecasting the Mawar region.
@@
During the 11-time-step (2.75 days) forecasting period, within the forecasting area (longitude 72°W to 30°W, latitude 0°N to 30°N), the MAE of the OCN model (0.17 m) was lower than that of the WW3 model (0.26 m). In terms of ACC, the minimum value for the OCN model (0.79) was slightly lower than that of the WW3 model (0.82). Regarding RMSE, the maximum value for the OCN model (0.03 m) was similar to that of the WW3 model (0.03 m). However, if we focus on the first 4 time steps (before the WW3 model updates its forecasting results), the minimum ACC value for the OCN model (0.98) was higher than that of the WW3 model (0.96), and the maximum RMSE value for the OCN model (0.02 m) was lower than that of the WW3 model (0.03 m). Based on these results, we can conclude that the short-term forecasting performance of the OCN model is comparable to that of traditional forecasting models.
@@ 3.3 Analysis of the Model’s Physical Properties
Figure 7: Verification experiment of the OCN model’s wave generation process under zero initial conditions. a) and c) show the global distribution of the absolute error of the OCN model at time step 1 and time step 63, respectively. It can be seen from the figures that at time step 63, the error of the OCN model is small in most sea areas, indicating that the model can well simulate the wave generation process. b) and d) show the ACC and RMSE curves of the model prediction results with respect to the forecast time. As the forecast time increases, the ACC gradually improves and stabilizes around 0.85, while the RMSE gradually decreases and stabilizes around 0.04m, further confirming the excellent performance of the OCN model in simulating the wave generation process. Overall, this experiment demonstrates that the OCN model can reproduce the wave generation process very well with high physical rationality.
During the 63-time-step running process, the model gradually self-corrected, and the final results tended to stabilize. The ACC value increased from -0.08 to around 0.85, while the RMSE decreased from 0.15 m to around 0.04 m. These trends demonstrate that the deep learning model exhibits similar characteristics to traditional physical models in terms of self-correction and convergence to a stable state. Specifically, the ACC value, which represents the correlation between the simulated and observed wave fields, started from a negative value (-0.08) due to the initially zero wave field. As the model ran, the ACC value gradually increased, indicating that the simulated wave field became increasingly consistent with the observed wave field. Similarly, the RMSE, which measures the average magnitude of the errors between the simulated and observed wave fields, started from a relatively high value (0.15 m) and decreased over time, indicating that the model’s accuracy improved as it self-corrected. These results suggest that the deep learning model, like traditional physical models, can capture the intrinsic relationships between wind and waves and simulate the dynamic processes of wave generation, propagation, and dissipation. This similarity in behavior supports the conclusion that the deep learning model has learned the underlying physical mechanisms governing wind-wave interactions.
@@
Figure 8: Figure shows the simulation performance of the OCN model during a complete typhoon process. a) and d) show the 10-meter wind speed distribution in the early (time step 1) and late (time step 95) stages of the simulation, respectively, presenting the entire process of typhoon generation, development, and dissipation. b) and e) show the spatial distribution of the absolute error of the OCN model at time steps 1 and 95, respectively. Comparing the two figures, it can be found that as the simulation time progresses, the error of the OCN model decreases significantly. Even in high wind speed areas such as the typhoon center, the model can provide relatively accurate forecasts in the later stage, indicating that the model can well simulate the wave evolution during the typhoon process. c) and f) show the ACC and RMSE curves of the model prediction results with respect to the forecast time. It can be seen that as the forecast time increases, the prediction skill of the model continuously improves, with ACC stabilizing at a high value of around 0.96 and RMSE decreasing to a low value of around 0.05 meters, quantitatively confirming the excellent performance of the OCN model in typhoon wave simulation.
During the 95-time-step simulation process, the ACC for the entire typhoon region increased from -0.13 to around 0.96, while the RMSE decreased from 0.18 m to around 0.05 m. The increasing ACC indicates that the simulated wave field became increasingly consistent with the observed wave field, while the decreasing RMSE suggests that the model’s accuracy improved over time. From the last time step, it can be observed that the errors in most areas were within 1 m, with the main error areas located near the typhoon. These results demonstrate that even under the influence of extreme wind fields, the model exhibits a convergence process similar to that of physical models.
@@
Figure 9: Simulation performance of the OCN model during Hurricane Lee. a) and d) show the 10-meter wind speed distribution at simulation time steps 1 and 55, respectively. b) and e) show the spatial distribution of the absolute error of the OCN model at time steps 1 and 55, respectively, indicating that the model error decreases significantly as the simulation progresses. c) and f) show the ACC and RMSE curves of the model prediction results with respect to the forecast time, where ACC eventually stabilizes around 0.97 and RMSE decreases to around 0.04 meters, confirming the excellent performance of the OCN model in the wave simulation of Hurricane Lee.
During the 55-time-step simulation process, the ACC for the entire hurricane region increased from 0.52 to around 0.97, while the RMSE decreased from 0.21 m to around 0.04 m. Similar to the case of Mawar, these trends indicate that the model’s performance improved over time, with the simulated wave field becoming increasingly consistent with the observed wave field. From the last time step, it can be observed that, similar to the situation with Mawar, the errors in most areas were within 1 m, with the main error areas located near the hurricane. These results further confirm that even under the influence of extreme wind fields, the model exhibits a convergence process similar to that of physical models. From the above results, it can be concluded that, even in extreme situations, the model demonstrates a convergence process similar to that of physical models. We can consider that the model exhibits a physical model-like convergence effect even under extreme conditions.

### v3

@@ 3.1 Performance Evaluation on ERA5 Dataset
During the 59 time steps (14.75 days) of the forecasting process, the model’s mean absolute error(MAE) is 0.08 m, with a minimum ACC of 0.82 and a maximum RMSE of 0.45 m. Considering the evaluation curves, the model performs well in forecasting the wave field. It is worth noting that although the model’s error decreases over time, it does not continue to decline like other forecasting models. Instead, it stabilizes at around 0.82 after the initial decrease. Based on this, we hypothesize that the model has learned the physical properties of waves.We will present relevant experiments in Section 3.3.
@@
During the 35-time-step (8.75 days) forecasting period, within the project area (longitude 120°E to 170°E, latitude 0°N to 50°N), the MAE was 0.24 m, the minimum ACC was 0.98, and the maximum RMSE was 0.32 m. Considering the evaluation curves, the model’s forecasting performance during Typhoon Goni can be regarded as satisfactory.
@@
During the 23-time-step (5.75 days) forecasting period, within the forecasting area (longitude 100°W to 60°W, latitude 0°N to 20°N), the MAE was 0.21 m, the minimum ACC was 0.84, and the maximum RMSE was 0.33 m. Considering the evaluation curves, the model’s forecasting performance in the hurricane region is also highly satisfactory.
@@ 3.2 Performance Comparison against Traditional Forecasting Models
During the first 7 time steps (1.75 days) of the forecasting process, the MAE of the OCN model was 0.14, which was lower than the MAE of the WW3 model (0.30). In terms of ACC, the minimum value for the OCN model was 0.99, higher than the 0.92 of the WW3 model. Regarding RMSE, the maximum value for the OCN model was 0.36, lower than the 1.26 of the WW3 model. Considering the evaluation curves, we can conclude that the short-term forecasting performance of the OCN model under normal conditions is better than that of the WW3 model.
@@
During the 11-time-step (2.75 days) forecasting period, within the forecasting area (longitude 120°E to 162°E, latitude 0°N to 30°N), the MAE of the OCN model (0.16 m) was lower than that of the WW3 model (0.26 m). In terms of ACC, the minimum value for the OCN model (0.98) was higher than that of the WW3 model (0.97). Regarding RMSE, the maximum value for the OCN model (0.51 m) was lower than that of the WW3 model (0.71 m). Considering the evaluation curves, we can conclude that the OCN model performed better than the WW3 model in forecasting the Mawar region.
@@
During the 11-time-step (2.75 days) forecasting period, within the forecasting area (longitude 72°W to 30°W, latitude 0°N to 30°N), the MAE of the OCN model (0.17 m) was lower than that of the WW3 model (0.26 m). In terms of ACC, the minimum value for the OCN model (0.79) was slightly lower than that of the WW3 model (0.82). Regarding RMSE, the maximum value for the OCN model (0.22 m) was similar to that of the WW3 model (0.21 m). However, if we focus on the first 4 time steps (before the WW3 model updates its forecasting results), the minimum ACC value for the OCN model (0.97) was higher than that of the WW3 model (0.91), and the maximum RMSE value for the OCN model (0.13 m) was lower than that of the WW3 model (0.21 m). Based on these results, we can conclude that the short-term forecasting performance of the OCN model is comparable to that of traditional forecasting models.
@@ 3.3 Analysis of the Model’s Physical Properties
Figure 7: Verification experiment of the OCN model’s wave generation process under zero initial conditions. a) and c) show the global distribution of the absolute error of the OCN model at time step 1 and time step 63, respectively. It can be seen from the figures that at time step 63, the error of the OCN model is small in most sea areas, indicating that the model can well simulate the wave generation process. b) and d) show the ACC and RMSE curves of the model prediction results with respect to the forecast time. As the forecast time increases, the ACC gradually improves and stabilizes around 0.85, while the RMSE gradually decreases and stabilizes around 0.55m, further confirming the excellent performance of the OCN model in simulating the wave generation process. Overall, this experiment demonstrates that the OCN model can reproduce the wave generation process very well with high physical rationality.
During the 63-time-step running process, the model gradually self-corrected, and the final results tended to stabilize. The ACC value increased from -0.08 to around 0.85, while the RMSE decreased from 2.36 m to around 0.55 m. These trends demonstrate that the deep learning model exhibits similar characteristics to traditional physical models in terms of self-correction and convergence to a stable state. Specifically, the ACC value, which represents the correlation between the simulated and observed wave fields, started from a negative value (-0.08) due to the initially zero wave field. As the model ran, the ACC value gradually increased, indicating that the simulated wave field became increasingly consistent with the observed wave field. Similarly, the RMSE, which measures the average magnitude of the errors between the simulated and observed wave fields, started from a relatively high value (2.36 m) and decreased over time, indicating that the model’s accuracy improved as it self-corrected. These results suggest that the deep learning model, like traditional physical models, can capture the intrinsic relationships between wind and waves and simulate the dynamic processes of wave generation, propagation, and dissipation. This similarity in behavior supports the conclusion that the deep learning model has learned the underlying physical mechanisms governing wind-wave interactions.
@@
Figure 8: Figure shows the simulation performance of the OCN model during a complete typhoon process. a) and d) show the 10-meter wind speed distribution in the early (time step 1) and late (time step 95) stages of the simulation, respectively, presenting the entire process of typhoon generation, development, and dissipation. b) and e) show the spatial distribution of the absolute error of the OCN model at time steps 1 and 95, respectively. Comparing the two figures, it can be found that as the simulation time progresses, the error of the OCN model decreases significantly. Even in high wind speed areas such as the typhoon center, the model can provide relatively accurate forecasts in the later stage, indicating that the model can well simulate the wave evolution during the typhoon process. c) and f) show the ACC and RMSE curves of the model prediction results with respect to the forecast time. It can be seen that as the forecast time increases, the prediction skill of the model continuously improves, with ACC stabilizing at a high value of around 0.96 and RMSE decreasing to a low value of around 0.26 meters, quantitatively confirming the excellent performance of the OCN model in typhoon wave simulation.
During the 95-time-step simulation process, the ACC for the entire typhoon region increased from -0.13 to around 0.96, while the RMSE decreased from 1.38 m to around 0.36 m. The increasing ACC indicates that the simulated wave field became increasingly consistent with the observed wave field, while the decreasing RMSE suggests that the model’s accuracy improved over time. From the last time step, it can be observed that the errors in most areas were within 1 m, with the main error areas located near the typhoon. These results demonstrate that even under the influence of extreme wind fields, the model exhibits a convergence process similar to that of physical models.
@@
Figure 9: Simulation performance of the OCN model during Hurricane Lee. a) and d) show the 10-meter wind speed distribution at simulation time steps 1 and 55, respectively. b) and e) show the spatial distribution of the absolute error of the OCN model at time steps 1 and 55, respectively, indicating that the model error decreases significantly as the simulation progresses. c) and f) show the ACC and RMSE curves of the model prediction results with respect to the forecast time, where ACC eventually stabilizes around 0.97 and RMSE decreases to around 0.26 meters, confirming the excellent performance of the OCN model in the wave simulation of Hurricane Lee.
During the 55-time-step simulation process, the ACC for the entire hurricane region increased from 0.52 to around 0.97, while the RMSE decreased from 1.78 m to around 0.26 m. Similar to the case of Mawar, these trends indicate that the model’s performance improved over time, with the simulated wave field becoming increasingly consistent with the observed wave field. From the last time step, it can be observed that, similar to the situation with Mawar, the errors in most areas were within 1 m, with the main error areas located near the hurricane. These results further confirm that even under the influence of extreme wind fields, the model exhibits a convergence process similar to that of physical models. From the above results, it can be concluded that, even in extreme situations, the model demonstrates a convergence process similar to that of physical models. We can consider that the model exhibits a physical model-like convergence effect even under extreme conditions.